Find the area bounded by y=cosx, the x-axis and between x=−π2 to x=π2. Draw a sketch for yourself to help you find the integral.
1
Expert's answer
2020-05-25T21:00:32-0400

"\int_{-\pi/2}^{\pi/2}\cos xdx=2\int_{0}^{\pi/2}\cos xdx=2\sin x|_0^{\pi/2}=2"